OpenStudy (anonymous):

The length of a rectangle is 7 1/2 in. and the width is 2 3/4in. What is the ratio, using whole numbers, of the length to the width?

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@louiseopura

OpenStudy (aravindg):

Just take the ratio?

OpenStudy (anonymous):

I don't under stnd how to

OpenStudy (aravindg):

First convert the mixed fractions to proper fractions then simply take ratio.

OpenStudy (anonymous):

how do you convert.... I really do not understand that's why I am asking this question sir

OpenStudy (aravindg):

7 1/2=(7*2+1)/2=15/2 Similarly do for the other

OpenStudy (anonymous):

that still confuse me im more confused now then eva

OpenStudy (aravindg):